OpenStudy (anonymous):

400m run. I'm faster than my partner by 4sec/ 1m. My partner finished the run in 50seconds. how do I create an equation and solve for this ?

OpenStudy (anonymous):

400/50=8m/sec.

OpenStudy (anonymous):

can you show your method ?

OpenStudy (anonymous):

so yur 8.25 so tim=400/8.25

OpenStudy (anonymous):

Okay Rubie why don't we start by calculating the speed of your partner. Now you will have to calculate it in sec/m unless that was a typo earlier. I will let you do the calculation and step in if necessary.

OpenStudy (anonymous):

vipul if you are wondering why I said to give it in sec/m it is because that is the way they gave the realtionship between the runners. I may not reply for a few minutes here I need to do something quickly.

OpenStudy (anonymous):

ohh okai so 400/50 ?

OpenStudy (anonymous):

no track your units. the numbers you gave me up there would render m/s the relationship they gave us was in s/m . you need to get those units

OpenStudy (anonymous):

nvm.... i'll just leave this question out then o.0

OpenStudy (anonymous):

50/400 is the correct first step giving you a speed of 0.125 s/m

OpenStudy (anonymous):

Unfortunately I think you have a typo here as if you are faster than your partner by 4 s/m then your speed is 0.125-4 = -3.75 s/m which is impossible

OpenStudy (anonymous):

\[\frac{ s }{ m }\] or \[\frac{ m }{ s }\] which is the unit your relationship is given in?